# Generative Modeling of Neural Dynamics via Latent Stochastic Differential Equations

### v1

@@ 4.1 Simulated data
| Coupled Oscillator SDE | 526 | 0.238 ± 0.008 | 0.86 ± 0.018 | 0.22 ± 0.011 | 0.67 ± 0.025 | 0.314 ± 0.01 |
4.2 Empirical data
4.2.1 Datasets & baselines
@@
Perturbed reach task In this experiment, a monkey is presented with a target location and tasked with moving the cursor from the center to the target location by controlling a manipulandum (Chowdhury et al., 2020) (Figure 3a). On one portion of the trials, the monkey’s arm was perturbed before the reach by applying resisting forces on the manipulandum. Neural activity was recorded from Brodmann’s area 2 of the somatosensory cortex. Here, we trained a model to predict stimulus-evoked neural responses in area 2 as well as the hand position of the monkey. The input to the model are the target location, as well as the perturbing force and torques on the manipulandum.
@@
Visual decision making In this task, a mouse is presented with images of different contrasts and the mouse is tasked with turning a wheel to bring the image with higher contrast to the center of the screen when a go cue is randomly presented (Steinmetz et al., 2019) (Figure 3d). Neural activity is recorded in multiple regions across the brain using Neuropixels probes. Here, we trained the model to predict stimulus-invoked neural responses across several brain regions simultaneously as well as predict the wheel velocity. The input to the model are the contrast levels, the timing of their presentation, and the timing of the go cue.
@@
Delayed reach task The task is a delayed center-out reach task with barriers, resulting in a variety of straight and curved trajectories (Mante et al., 2013) (Figure 3g). Neural activity was recorded from the dorsal premotor (PmD) and primary motor (M1) cortices, and cursor, monkey gaze position, and monkey hand position and velocity are also provided. Here we train a model to predict neural responses during movement in both areas given the preparatory activity before the go cue. No stimulus information is provided to the model during training or prediction.
@@
We compare different latent variable models across the three datasets. To enable a fair compariosn, we fix the encoder and decoder architectures across all models while varying the latent sequential model. That is, we compare the performance of RNN, GRU, LSTM, neural ODE, and coupled oscillator ODE models with that of our proposed coupled oscillator SDE and a neural SDE model. All the baseline models had a probabilistic initial state and were trained via variational inference. Note that this formulation enables generalizing the baselines to several previous models proposed in literature (Pandarinath et al., 2018; Hurwitz et al., 2021b; Kim et al., 2021; Versteeg et al., 2023) (See Appendix D for details on the model architecture).
4.2.2 Empirical results
@@
Table 1 presents 5-fold cross-validation results across the three datasets, evaluating both neural activity prediction (in bits per spike) and behavioral performance (via the coefficient of determination, ). First, we observe that our stochastic models (Neural SDE and CO-SDE) on average consistently outperform their deterministic counterparts. This suggests that incorporating stochasticity into the latent dynamics is crucial for capturing the variability inherent in neural and behavioral data and that a stochastic initial state might not be sufficient to capture uncertainty in the dynamics.
@@
Second, we observe that model complexity, as measured by parameter count, does not correlate with performance. The coupled oscillator models achieve competitive or superior performance while using only 465-526 parameters, compared to 33,280 parameters for the LSTM model. Note that while this number does not account for the number of parameters required for modeling the approximate posterior SDE (3528), this remains a significant improvement in parameter efficiency for the generative model. This suggests that incorporating appropriate inductive biases through the coupled oscillator architecture enables more efficient modeling of the underlying dynamics. This also suggests that standard deep learning models may be too over-parameterized to learn the underlying dynamical system during simple behavioral tasks.
Beyond quantitative results, oscillator-based models provide a direct way to interpret latent neural dynamics. We visualize the value of the inferred frequencies of the model after training as well as the value of the coupling strength in response to sample stimuli in Figure 3. In the perturbed reach task, where we have data from one somatosensory area, we observe that the connectivity strength increases over time (Figure 3c). This is also observed in the delayed reach task where we have recordings from two regions in the motor in the motor cortex (Figure 3k). In contrast, on the visual decision making task where we have recordings from multiple regions across the brain the coupling strength decreases over time (Figure 3f). This decrease is marked via two significant drops at two timepoints which mark the presentation of the stimulus and the go cue, respectively. Additionally, we observe that on average, the inferred natural frequency of the oscillators differ between brain regions. Notably, the average frequency of oscillators corresponding to M1 is higher than the average frequency of oscillators corresponding to PMd during the delayed reach task (Figure 3k), which may be linked to distinct functional roles related to action preparation and execution (Thura et al., 2022).
5 Discussion
@@
First, explicitly modeling probabilistic dynamics through SDEs consistently improves model performance compared to deterministic alternatives, even when the latter incorporate uncertainty through probabilistic initial conditions. This suggests that neural variability stems not just from uncertain initial states but from inherently stochastic dynamics (Laing & Lord, 2009). The superior performance of SDE-based models in high-noise regimes as shown in the simulations results further supports this conclusion. Additional simulation results also show that using stochastic dynamics can improve model fit while relying on lower-dimensional latent states; an observation that is also found in stochastic low-rank RNNs (Pals et al., 2024).
@@
Second, our hybrid coupled oscillator model achieves competitive or superior performance while requiring an order of magnitude fewer parameters than standard deep learning approaches. This parameter efficiency suggests that incorporating appropriate inductive biases through mechanistic components can effectively constrain the model space (Rackauckas et al., 2020; ElGazzar & van Gerven, 2024). The success of coupled oscillators in particular hints at their potential as a natural description language for neural population dynamics, consistent with existing theoretical work on neural oscillations (Buzsaki & Draguhn, 2004; Breakspear et al., 2010; Churchland et al., 2012; Bick et al., 2020). This also supports the potential of coupled oscillators as an expressive and efficient tool to model arbitrary time-series data, as seen in emerging machine learning research (Rusch & Mishra, 2020; Effenberger et al., 2022; Rusch & Rus, 2024).
@@
With that said, there are a number of limitations to the proposed approach which merits consideration. First, the framework relies on variational inference for training the model parameters. This could be suboptimal as variational inference is known to provide overconfident uncertainty estimates and is prone to several optimization challenges such as posterior collapse, especially when coupled with powerful decoders (Blei et al., 2017). Additionally, training the model parameters currently requires back-propagation through an SDE solver, which can be computationally expensive in terms of both time and memory requirements. Potential solutions could lie in emerging simulation-free approaches for training (neural) stochastic differential equations (Course & Nair, 2024; Zhang et al., 2024). Future work could also explore comparing different dynamical systems beyond coupled oscillators and focus on extending the current approach to model neural data from several subjects as well as other recording modalities.
@@ Acknowledgements
This work is supported via the Dutch Brain Interface Initiative (DBI2) with project number 024.005.022 of the research programme Gravitation which is (partly) financed by the Dutch Research Council (NWO).
Appendix
Appendix A Related Work
Latent variable models in neuroscience
Our work builds on previous latent variable models which have been proposed to infer low-dimensional latent dynamics from single-trial data (Sahani, 1999; Yu et al., 2008; Macke et al., 2011; Wu et al., 2017; Pandarinath et al., 2018; Hurwitz et al., 2021b; Schimel et al., 2021; Kim et al., 2021). Most related to our work, is recent work by Hu et al. (2024) which leverage SDEs to represent latent neural dynamics. In their work, the authors developed a novel Gaussian process kernel function that defines a smooth, locally linear prior on dynamics, enabling it to act as an easily interpetable recurrent linear switching dynamical systems while providing posterior uncertainty estimates. A key advantage of our employed approach is the scalability and flexibility which is limited by the use of Gaussian processes. Relevant recent work also by Pals et al. (2024) extend low-rank RNNs to a stochastic setting with additive process noise. The authors use a specific form of the RNN which enables finding fixed-points effectively after training.
Scalable inference of latent SDEs
Inferring the state and learning the parameters of a latent SDE from (high-dimensional) noisy observations is a challenging problem which appears through out science and engineering. With the advent of neural differential equations (Chen et al., 2018; Kidger, 2022), there has been a growing interest in developing scalable methods for fitting latent neural SDEs to data. Seminal work by Tzen & Raginsky (2019a, b) established the theoretical foundation for training generative Neural SDEs via variational inference through the lens of stochastic control. Subsequent research has focused on addressing key computational challenges, including improved memory efficiency and extension to the time domain (Li et al., 2020), enhanced numerical stability (Zeng et al., 2023; Oh et al., 2024), and computational efficiency (Course & Nair, 2024). Our work builds upon the formulation by Li et al. (2020), extending it to controlled and multi-modal settings in neuroscience. This specific formulation enables preserving the expressivity of the generative SDE through state- and input-dependent diffusion on arbitrary spaces.
Modeling neural dynamics with coupled oscillators
Coupled oscillator models have a rich history in theoretical neuroscience, dating back to seminal work by Kuramoto (1975) and Wilson & Cowan (1972). These models have proven particularly effective in capturing rhythmic neural activity and investigating synchronization phenomena in neural circuits (Breakspear et al., 2010; Ermentrout & Terman, 2010). Of particular relevance to our work are modern approaches that combine oscillator models with machine learning techniques (Abrevaya et al., 2021, 2023; Bandyopadhyay et al., 2023) to learn dynamical systems that can reproduce complex neural trajectories while maintaining interpretable structure. Our approach builds upon these foundations by introducing a flexible SDE framework that can naturally capture oscillatory dynamics while accounting for process noise and external inputs. This bridges the gap between mechanistic oscillator models and data-driven approaches, allowing for both accurate trajectory prediction and meaningful scientific interpretation.
Appendix B Learning the parameters of the generative model and inferring latent paths
@@
Appendix C Implementation details
In the following, we summarize the general implementation details of our latent SDE. The models were trained to minimize ELBO on the training set by standard stochastic gradient backprogation through the computational graph. This includes the numerical solver. That is, the intermediate steps are saved in memory during the forward pass and used to update the parameters during the backward pass. While this scales to in terms of memory requirement where is the number of steps in the numerical solver compared to of the adjoint method, we found that this leads to more stable training in our experiments. We used an Euler-Maruyama solver for all our experiments. We found that setting a step size at 2 times the bin width of the spikes in our empirical experiments did not significantly influence the performance and helped improved the training speed and memory requirements.
Notably, we found that cyclic annealing the KL-divergence term in our ELBO significantly improved training stability. We follow the approach by Fu et al. (2019) and use a linear with 4 cycles with a linearly increasing value from 0 to 10 for half the cycle and then stays constant for the second half, until the cycle resets. In experiments which include reconstructing both neural and behavioral data, we use a constant weighting for each during training depending on the dataset. Our experiments were carried out on a NVIDIA A100-PCIE-40GB. Table 2 lists the main components of our latent SDE and their hyperparameters used across all empirical experiments.
@@ Appendix D Architecture
The observation encoder serves two purposes: (1) determining the parameters of the posterior distributions for the initial state, and (2) generating a lower-dimensional time-series representation (context) to condition the augmented SDE on the data. This is accomplished through two parallel components: the initialization network and the context network. The initialization network processes the first 30The context network processes the complete time-series of observations through a 2-layer LSTM, producing a time-series of length 64. This context vector undergoes linear interpolation to create a continuous-time representation for the augmented SDE.
@@
The generative SDE defines our dynamics by mapping the current state and encoded stimulus to the change in the next state. The drift and diffusion structure is flexible and should incorporate prior knowledge when available. In this paper, we present a model combining coupled oscillators with neural networks to represent the SDE’s drift vector field. We also evaluate a neural SDE variant using a state size of 16 and a 2-layer MLP with hidden layer size 64 and tanh activation functions. The diffusion uses a sparse connection layer with non-zero diagonal parameters only, which is shared with the augmented SDE.

### v2

@@ Results
| CO-SDE | 526 | 0.238 ± 0.008 | 0.86 ± 0.018 | 0.22 ± 0.011 | 0.67 ± 0.025 | 0.314 ± 0.01 |
Datasets & baselines
@@
In this experiment, a monkey is presented with a target location and tasked with moving the cursor from the center to the target location by controlling a manipulandum [12] (Fig. 3a). On one portion of the trials, the monkey’s arm was perturbed before the reach by applying resisting forces on the manipulandum. Neural activity was recorded from Brodmann’s area 2 of the somatosensory cortex. Here, we trained a model to predict stimulus-evoked neural responses in area 2 as well as the hand position of the monkey. The input to the model are the target location, as well as the perturbing force and torques on the manipulandum. At test-time, we provide the first 50 ms of observations before stimulus presentation to the observation encoder to identify the initial condition.
Visual decision making
@@
In this task, a mouse is presented with images of different contrasts and the mouse is tasked with turning a wheel to bring the image with higher contrast to the center of the screen when a go cue is randomly presented [71] (Fig. 3d). Neural activity is recorded in multiple regions across the brain using Neuropixels probes. Here, we trained the model to predict stimulus-invoked neural responses across several brain regions simultaneously as well as predict the wheel velocity. The input to the model are the contrast levels, the timing of their presentation, and the timing of the go cue. At test-time, we provide the first 50 ms of observations before stimulus presentation to the observation encoder to identify the initial conditions.
Delayed reach task
@@
The task is a delayed center-out reach task with barriers, resulting in a variety of straight and curved trajectories [52] (Fig. 3g). Neural activity was recorded from the dorsal pre-motor (PmD) and primary motor (M1) cortices, and cursor, monkey gaze position, and monkey hand position and velocity are also provided. Here we train a model to predict neural responses during movement in both areas given the preparatory activity before the go cue. No stimulus information is provided to the model during training or prediction. At test-time, we provide the first 100 ms of neural observations before the go cue to the observation encoder to identify the initial conditions.
@@
We compare different latent variable models across the three datasets. To enable a fair comparison, we fix the encoder and decoder architectures across all models while varying the latent sequential model. That is, we compare the performance of RNN, GRU, LSTM, Neural ODE, and a coupled oscillator ODE (CO-ODE) models with that of our proposed CO-SDE and a Neural SDE model. We also include a linear SDE model which employs linear drift and diffusion functions. All the baseline models had a probabilistic initial state and were trained via variational inference. Note that this formulation enables generalizing the baselines to several previous models proposed in literature [57, 37, 43, 80]. We also compare against a recently proposed Gaussian process switching linear dynamical system (GP-SLDS) [2] (See Supplementary Information for details on the baselines).
Empirical results
@@
Table 1 presents 5-fold cross-validation results across the three datasets, evaluating both neural activity prediction (in bits per spike) and behavioral performance (via the coefficient of determination, ) (see Supplementary Information for further details on the evaluation metrics). First, we observe that our stochastic models (Neural SDE and CO-SDE) on average consistently outperform their deterministic counterparts. This suggests that incorporating stochasticity into the latent dynamics is crucial for capturing the variability inherent in neural and behavioral data and that a stochastic initial state might not be sufficient to capture uncertainty in the dynamics.
Second, we find that non-linear latent dynamics are crucial for accurate joint prediction of neural and behavioral responses. This is suggested by the performance of the linear SDE baseline which performs competitively to non-linear baselines only on predicting behavioral responses on the perturbed reach task. However, its performance is significantly lower for predicting neural responses across all three tasks. This suggests that while linear dynamics may suffice for capturing certain behavioral regularities, nonlinear models are necessary for accurate neural population modeling at the level of single-neuron.
@@
Third, we observe that model complexity, as measured by parameter count, does not correlate with performance. The coupled oscillator models achieve competitive or superior performance while using only 465-526 parameters, compared to 33,280 parameters for the LSTM model. Note that while this number does not account for the number of parameters required for modeling the approximate posterior SDE (3528), this remains a significant improvement in parameter efficiency for the generative model. This suggests that incorporating appropriate inductive biases through the coupled oscillator architecture enables more efficient modeling of the underlying dynamics. This also suggests that standard deep learning models may be too over-parameterized to learn the underlying dynamical system during simple behavioral tasks.
Beyond quantitative performance, oscillator-based models offer interpretable parameterizations that connect to established neuroscience concepts. We visualize the inferred natural frequencies and coupling strengths learned by the model in Fig. 3. In both motor tasks —the perturbed reach task with recordings from somatosensory cortex and the delayed reach task with recordings from two motor cortical areas- we observe that coupling strength increases over time during movement preparation and execution (Fig.3c,k). This pattern is consistent with theories of neural communication through enhanced synchronization [26], where increasing oscillatory coupling may facilitate information integration across neural populations during coordinated motor output. In contrast, the visual decision-making task shows a markedly different pattern (Fig.3f). Here, coupling strength decreases over time, with two distinct drops coinciding with the stimulus presentation and go cue. This temporal modulation may reflect dynamic routing of information[26], where coupling is selectively reduced to gate irrelevant signals or reallocate processing resources as task demands shift from sensory encoding to decision formation and action selection. We also observe systematic differences in natural frequencies across brain regions. Notably, oscillators corresponding to M1 exhibit higher average frequencies than those in PMd during the delayed reach task (Fig.3h). This frequency gradient is consistent with the hierarchical organization of motor cortex, where M1 operates at faster timescales than PMd, which is more involved in movement planning [73, 13]. Such frequency differences may reflect distinct computational roles: higher frequencies in M1 may support rapid sensorimotor transformations required for real-time motor control, while lower frequencies in PMd may be suited for integrating information over longer timescales during motor preparation. Important note on interpretation: While these learned parameters show relative consistency across model initializations and align with neuroscience literature, we emphasize that our method does not provide theoretical guarantees for parameter identifiability—a well-known challenge in nonlinear dynamical systems [54]. Therefore, we present these findings as hypothesis-generating observations that suggest our model captures functionally meaningful structure, rather than as definitive claims about the underlying neural mechanisms. The value of these parameterizations lies in providing neuroscience-motivated, interpretable descriptions of population dynamics that can guide future experimental investigations.
Discussion
@@
First, explicitly modeling probabilistic dynamics through SDEs consistently improves model performance compared to deterministic alternatives, even when the latter incorporate uncertainty through probabilistic initial conditions. This suggests that neural variability stems not just from uncertain initial states but from inherently stochastic dynamics [48]. The superior performance of SDE-based models in high-noise regimes as shown in the simulations results further supports this conclusion. Additional simulation results also show that using stochastic dynamics can improve model fit while relying on lower-dimensional latent states; an observation that is also found in stochastic low-rank RNNs [56].
@@
Second, our hybrid coupled oscillator model achieves competitive or superior performance while requiring an order of magnitude fewer parameters than standard deep learning approaches. This parameter efficiency suggests that incorporating appropriate inductive biases through mechanistic components can effectively constrain the model space [61, 20]. The success of coupled oscillators in particular hints at their potential as a natural description language for neural population dynamics, consistent with existing theoretical work on neural oscillations [10, 8, 13, 6]. This also supports the potential of coupled oscillators as an expressive and efficient tool to model arbitrary time-series data, as seen in emerging machine learning research [66, 19, 67]. Future work could explore in depth the connection between features of inferred coupled oscillators (e.g. synchronization, phase-amplitude coupling, etc.) and observed neural and behavioral responses.
While we focus on coupled oscillator models motivated by motor cortex physiology, our framework readily accommodates diverse dynamical systems including Van der Pol oscillators [77] for modeling limit cycles with different relaxation properties, FitzHugh-Nagumo [24] models for excitable systems, attractor-based models for decision-making circuits [82], or predator-prey dynamics for competitive interactions [27]. The choice of dynamical model should be guided by domain knowledge and scientific hypotheses about the neural system. This flexibility enables researchers to test specific mechanistic hypotheses while maintaining the benefits of probabilistic inference and uncertainty quantification. Future work could systematically compare these alternative models on the same datasets to identify which dynamical structures best capture different neural computations. Alternatively, the symbolic expressions that define the latent dynamics could be inferred using black-box optimization techniques such as differential genetic programming [79].
@@
With that said, there are a number of limitations to the proposed framework which merits consideration. First, similar to most non-linear latent variable models in the literature, our approach does not guarantee parameter identifiability. Rather, our approach enables interpreting the system dynamics as a function of the behavior of the proposed system dynamics. Further work is required to ensure identifiability of the parameters of the latent dynamical system to enable robust comparison of different models [17, 32]. Here, advances in epistemic AI that focus on quantifying parameter uncertainty are of particular relevance [51]. Another limitation is that our framework relies on variational inference for training the model parameters. This could be suboptimal as variational inference is known to provide overconfident uncertainty estimates and is prone to several optimization challenges such as posterior collapse, especially when coupled with powerful decoders [7]. That is, the model can rely on expressive decoders to model the data, rendering latent variables redundant. A potential solution is to utilize invertible decoders [89, 80] to guarantee that the model relies on latent dynamics to model neural and behavioral observations. Furthermore, training the model parameters currently requires back-propagation through an SDE solver, which can be computationally expensive in terms of both time and memory requirements. Potential solutions could lie in emerging simulation-free approaches for training (neural) stochastic differential equations [14, 88, 5].
In this work, we provided a first demonstration of the usefulness of developing biophysics-inspired latent SDE models. Future work could explore comparing different dynamical systems beyond coupled oscillators and focus on extending the current approach to model neural data from several subjects as well as other recording modalities. The use of these models for neural system identification may provide new insights into neural information processing. Furthermore, their interpretability and expressiveness may pave the way for new applications in clinical neuroscience.
@@ Acknowledgements
This publication is part of the project Dutch Brain Interface Initiative (DBI2) with project number 024.005.022 of the research programme Gravitation which is (partly) financed by the Dutch Research Council (NWO).
Supplementary Information
1 Learning the parameters of the generative model and inferring latent paths
@@
2 Implementation details
In the following, we summarize the general implementation details of our latent SDE. The models were trained to minimize ELBO on the training set by standard stochastic gradient backprogation through the computational graph. This includes the numerical solver. That is, the intermediate steps are saved in memory during the forward pass and used to update the parameters during the backward pass. While this scales to in terms of memory requirement where is the number of steps in the numerical solver compared to of the adjoint method, we found that this leads to more stable training in our experiments. We used an Euler-Maruyama solver for all our experiments. We found that setting a step size at 2 times the bin width of the spikes in our empirical experiments did not significantly influence the performance and helped improved the training speed and memory requirements. We ran a hyperparameter search for the latent states, which in our oscillatory models (CO-ODE, CO-SDE) represents the number of oscillators . We found that the optimal number of oscillators varies across tasks (typically between 8 and 16) for the evalutaed tasks. We also found that cyclic annealing the KL-divergence term in our ELBO significantly improved training stability. We use a linear with 4 cycles with a linearly increasing value from 0 to 10 for half the cycle and then stays constant for the second half, until the cycle resets. In experiments which include reconstructing both neural and behavioral data, we use a constant weighting for each during training depending on the dataset. Our experiments were carried out on a NVIDIA A100-PCIE-40GB. Table 2 lists the main components of our latent SDE and their hyperparameters used across all empirical experiments.
@@ 3 Architecture
The observation encoder serves two purposes: (1) determining the parameters of the posterior distributions for the initial state, and (2) generating a lower-dimensional time-series representation (context) to condition the augmented SDE on the data. This is accomplished through two parallel components: the initialization network and the context network. The initialization network processes the first 30% of the observations using a 2-layer LSTM with hidden size 64, applying it in reverse order. The network then transforms the output through two parallel affine transformations into the latent dimension size, with a soft-plus activation applied to the variance component. This generates the mean and squared variance of the posterior initial condition. The context network processes the complete time-series of observations through a 2-layer LSTM, producing a time-series of length 64. This context vector undergoes linear interpolation to create a continuous-time representation for the augmented SDE.
@@
The generative SDE defines our dynamics by mapping the current state and encoded stimulus to the change in the next state. The drift and diffusion structure is flexible and should incorporate prior knowledge when available. In this paper, we present a model combining coupled oscillators with neural networks to represent the SDE’s drift vector field. We also evaluate a neural SDE variant with a 2-layer MLP with hidden layer size 64 and tanh activation functions. The diffusion uses a sparse connection layer with non-zero diagonal parameters only, which is shared with the augmented SDE.
@@
4 Baseline Model Specifications
To ensure a fair comparison across all models, we fix the encoder and decoder architectures while varying only the latent sequential dynamics model. This section provides comprehensive architectural details and parameter counts for each model variant.
Training Configuration
All models (except gpSLDS) were trained using the hyperparameters specified in Table 2. Key training details include:
All models trained via variational inference with ELBO objective
Probabilistic initial state: where and are produced by the shared observation encoder
For SDE models: Euler-Maruyama solver with step size
For ODE models: Euler solver with same step size
For RNN/GRU/LSTM: Standard discrete-time updates at observation time points
Cyclic KL annealing applied to all variational models
Implementation Notes
Neural ODE & Neural SDE
These models use flexible neural network parameterizations for the drift (and diffusion) functions. The MLP architecture allows the model to learn arbitrary smooth dynamics while maintaining computational tractability through modern ODE/SDE solvers.
Coupled Oscillator Models (CO-ODE & CO-SDE)
The coupled oscillator models incorporate mechanistic priors about neural oscillations. Each oscillator evolves according to:
where the coupling strengths are modulated by the stimulus through a small neural network. This inductive bias significantly reduces the parameter count while maintaining expressiveness.
Recurrent Models (RNN, GRU, LSTM)
These discrete-time models are evaluated at observation time points. To ensure fair comparison with continuous-time models, we:
Sample hidden states at the same time points as ODE/SDE models
Use the same batch size and training duration
Apply the same variational inference framework with probabilistic initial states
Fair Comparison Justification
By fixing encoder and decoder architectures, we isolate the impact of different latent dynamics models on performance. Any differences in results can be attributed solely to how well each model captures the temporal dynamics of neural and behavioral data. The parameter counts in Table 1 of the main text reflect only the latent dynamics parameters, as the encoder/decoder parameters (shared across all models) do not contribute to performance differences.
Evaluation Metrics
Bits per Spike
We assess the quality of neural response predictions using the bits per spike (bps) metric, an information-theoretic quantity that measures the gain in predictive log-likelihood of a model over a time-homogeneous Poisson baseline, normalized by the number of spikes. The metric is grounded in the Poisson log-likelihood and has been widely adopted in recent neural latent variable modeling literature.
Let denote the observed spike count for neuron at time bin , and let denote the firing rate predicted by the model. Let denote the empirical mean firing rate of neuron over the evaluation data. The bits per spike is computed as:
This expression reflects the difference in log-likelihood (in bits) between the model and a baseline that predicts a constant firing rate for each neuron. Positive values indicate better predictive performance than the baseline; zero corresponds to parity; negative values indicate worse-than-baseline predictions.
Our use of this metric is inspired by [1], who introduced bits per spike as a normalized log-likelihood score in the context of co-smoothing, wherein the goal is to predict the activity of held-out neurons conditioned on held-in neurons on the same trial. In contrast, we apply this metric to all recorded neurons on entirely leave-out trials that are not seen during training. This choice evaluates the model’s ability to generalize temporally, rather than across neurons, and aligns with our objective of modeling neural dynamics over continuous time.
Coefficient of Determination ()
To evaluate the model’s ability to predict behavioral measurements (e.g., kinematic variables or choice-related signals), we use the coefficient of determination, , under a Gaussian readout from the inferred latent state. This metric measures the proportion of variance in the observed behavior explained by the model’s predictions.
Let be the observed behavioral measurement at time , and the corresponding model prediction. The score is defined as:
where is the empirical mean of over the evaluation set. An score of 1 indicates perfect prediction, 0 corresponds to performance equal to predicting the mean, and negative values reflect worse-than-mean predictions. This metric provides an interpretable measure of behavioral predictability from latent neural dynamics.
Supplementary References